OpenStudy (anonymous):

HELP PLEASE!!(:

OpenStudy (anonymous):

Karen wrote the following steps to simplify the expression: 4 – (2x + 3x – 1).

OpenStudy (anonymous):

Which step of justification is incorrect, and what should the justification for that step be to simplify the expression? Step 2; Commutative Property of Addition Step 3; Associative Property of Multiplication Step 4; Distributive Property Step 1; Multiplication Property of Equality

OpenStudy (anonymous):

(Well you have to solve the problem in order to know what she did wrong) Look at the equation. From what you have learned, what do you have to do first?

OpenStudy (anonymous):

actually scratch that, lets go through the steps of what she did

OpenStudy (anonymous):

are you able to see the attachment?

OpenStudy (anonymous):

I am able to see it

OpenStudy (anonymous):

What was the first step she did? It was the distributive property, do you know how to do that?

OpenStudy (anonymous):

I am now seeing these messgaes cuz openstudy is SLOW

OpenStudy (anonymous):

nd yes I know how to do dist. property

OpenStudy (anonymous):

Okay its fine, so did this girl do it correctly??

OpenStudy (anonymous):

shouldn't it be 4 -( 5x-1)

OpenStudy (anonymous):

No, thats not correct for this step. She has the distributive property correct

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

So go to the next step. Does she have that correctly?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

well I don't think so

OpenStudy (anonymous):

correct she has that one wrong :) and thats your answer

OpenStudy (anonymous):

yayyy! thank you(:

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

can you help me with another please?

OpenStudy (anonymous):

I can try, this one almost had me stuck haha. I'm not the best at math but I'll give it a shot

OpenStudy (anonymous):

okay thanx!

OpenStudy (anonymous):

Given: 3x + 1 = 2 + 2x – 4 Prove: x = –3 Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2. Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above?

OpenStudy (anonymous):

I hate these-_- they are so hard

OpenStudy (anonymous):

@jewelgurl ??

OpenStudy (anonymous):

one moment (I'll get back to you in a second. I'm trying to multitask ahhaha)

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

@flykicks ... any ideas?

OpenStudy (anonymous):

Okay, I'm so sorry... i have absolutely no clue. :(